I am trying to help my daughter and am at a loss.Your help would be very much appreciated We are doing absolute value.Can you solve this equation |x-1| = 5x + 10
Answer by stanbon(75887)

|x-1| = 5x + 10
-------
Draw a number line.
Plot x = 1 on the line
----
If x < 1, x-1 is negative.
So |x-1| = 1-x
Solve: |x-1| = 5x+10
1-x = 5x+10
6x = -9
x = -3/2
=======================
If x > 1, x-1 is positive.
So |x-1| = x-1
Solve: x-1 = 5x+10
4x = 11
x = 11/4
=======================
Cheers,
Stan H.